Emma has had a puncture and the air is leaking from the tire. In addition, it is starting to rain. From the beginning, the tire held 5 dl of air with a pressure of 300 kPa. We can assume that the temperature is constant and that the average pressure during the leak is 165 kPa (the average pressure is obtained by drawing the graph using the ideal gas laws, but we have calculated it for you.

What volume does the gas get at normal air pressure?
1
Expert's answer
2021-03-14T19:17:11-0400
"(P_i-P_a)V_i=(P_f-P_a)V_f\\(300-101.325)5=(165-101.325)V_f\\V_f=15.6\ dL=1.56\ L"